OpenStudy (anonymous):

A car company tested a sports car on a road with different inclines. The test driver tested the car by driving a distance of x miles on a flat road, (x2 + 3) miles downhill, and (x − 7) miles uphill. Which simplified expression is equivalent to the total distance, in miles, for which the car was tested? 3x2 − 4 3x2 + 10 x2 + 2x − 4 x2 + 2x + 10

OpenStudy (michele_laino):

your problem asks for total distance d, which is: \[\Large d = x + \left( {{x^2} + 3} \right) + \left( {x - 7} \right) = ...\] please simplify